# InterQnet: A Heterogeneous Full-Stack Approach to Co-designing Scalable Quantum Networks

### v1

@@ 6.1.1 Simulating Optical Quantum Communications
We employ quantum trajectory theory [87]—which simulates open quantum system dynamics—to model the evolution of system statistics throughout the experiment. Kraus operators corresponding to the relevant noise models (e.g., amplitude damping, amplification, and depolarization) are applied to the matrix product states as they evolve through the system. These Kraus operators are applied probabilistically, with probabilities determined by the unnormalized weights of each trajectory [88]. Computing these probabilities for every iteration of a Monte Carlo experiment is computationally expensive and inefficient. Therefore, we developed a novel trajectory caching scheme. Since most quantum channels have certain trajectories that occur far more frequently than others, the required calculations for these common trajectories are performed only once, after which the cached results are reused for the majority of Monte Carlo samples.
6.1.2 Simulating Physical Qubits and Their Emission
@@
6.2 Entanglement witnessing in the logical plane of quantum networks
@@
6.3 Architectural Approaches to the Control and Service Plane of Scalable Quantum Networks
@@
In prior work [92] we proposed a theoretical framework that enables the continuous generation of entangled pairs between neighboring nodes. This approach can adapt to incoming client requests to reduce the time to serve future requests. In [92], however, the theoretical framework is evaluated only through small-scale, ad hoc, and naive time-slotted simulations. The challenge is to implement the theoretical framework into a comprehensive quantum network protocol and simulate the protocol in a discrete-event quantum network simulator at a large scale, including realistic physical parameters.
@@
The challenges are addressed in [93], where we designed and implemented the Adaptive, Continuous entanglement generation Protocol (ACP). ACP reduces the time to serve a client request by letting a node attempt to generate entangled pairs with neighboring nodes in the background continuously, regardless of incoming requests. The selection of neighbor nodes adapts to the request traffic by a feedback control algorithm. Additionally, entanglement distillation is used to enhance the fidelity of entangled pairs. We extended the SeQUeNCe quantum network simulator to support the simulation of ACP. In particular, we added a single-heralded entanglement generation protocol that accounts for pregenerated entangled pairs.
Figure 8: Simulation results for the 200-node autonomous system topology network.
We showed promising results when we evaluated the ACP in the SeQUeNCe simulator under various scales (2-node, 20-node, and 200-node networks) and realistic physical parameters (Table 1 in [93]). For instance, Fig. 8 shows the performance of ACP in the 200-node autonomous system topology network, comparing with on-demand-only (ODO) and uniform-continuous protocol (UCP). Fig. 8(a) shows the effectiveness of the continuous entanglement generation by UCP having a slightly lower time to serve than ODO. Fig. 8(a) also shows the effectiveness of the adaptiveness of ACP by having a 50+% lower time to serve than UCP. Adapting to incoming requests is a process that takes time, and explaining ACP’s red line gradually decreases. The sudden change in the request traffic pattern results in a spike for ACP’s time to serve (red line). Then, ACP will gradually adapt to the new request traffic pattern. Fig. 8(b) shows that ACP enhanced with entanglement distillation increases the fidelity of the end-to-end entangled pairs by as much as 0.05 compared with UCP and 0.04 compared with ODO, both without entanglement distillation.
6.4 Summary of Accomplishments and Open Challenges
The three-plane abstraction provides a unifying framework for reasoning about quantum network architectures. The infrastructure plane captures heterogeneous devices and links, the logical plane describes the generation, purification, and verification of entanglement, and the control/service plane coordinates these resources to support user applications. Our architectural analysis highlights that scalability depends not only on distance but also on node and user counts, heterogeneity, and entanglement distribution rates. These studies are designed to incorporate the error management methods of Section 5 as implementations mature, enabling evaluation of purification, distillation, and error correction in the context of realistic network protocols.
We have made significant progress in creating faithful and computationally efficient simulations of the infrastructure plane of quantum networks. With these tools, we discovered that visibility of entanglement swapping varies depending on measurement basis and source brightness, an effect arising from multipair emissions and realistic threshold detection. Our adaptive continuous entanglement generation protocol significantly improves the performance of scalable quantum networks, reducing the time to serve by up to 94% and improving fidelity by 0.05 in large-scale simulations. We are now scaling and optimizing this for hybrid links, larger networks, and centralized vs. decentralized control paradigms. Furthermore, we have made progress on SeQUeNCe integration, enhancing physical layer realism (see Section 4) and implementing protocol-level optimizations. Future work includes integrating devices and error correction models and comparing centralized vs. distributed control schemes in the control and service plane of quantum networks.
Collectively, these efforts build a scalable, realistic foundation for deploying and evaluating next-generation quantum networks. They provide the architectural context for InterQnet-Achieve demonstrations in ARQNET and for InterQnet-Scale simulations that explore qLDPC-based purification and large-scale orchestration. Section 7 builds on these results by describing how architectural protocols are realized in practice through integrated orchestration and control.
7 Integration
@@
Integration efforts on ARQNET bring the architectural concepts of Section 6 into practice, focusing on the coexistence of quantum and classical signals, stabilization of quantum channels under realistic fiber conditions, and orchestration of heterogeneous devices. These activities are essential to InterQnet-Achieve, where a three-node heterogeneous network will demonstrate entanglement distribution across superconducting, erbium, and ytterbium platforms, and they also inform InterQnet-Scale by identifying key constraints and performance trade-offs for realistic architectures.
7.1 Quantum-Classical Coexistence over Deployed Fiber
@@
We prototyped an on-demand entanglement distribution service over ARQNET, addressing the challenge of spontaneous Raman scattering (SpRS) induced by coexisting classical signals [29]. We developed an empirical model that quantitatively describes the relationship between the coincidences-to-accidentals ratio (CAR) and pump power of entangled photon sources in the presence of classical traffic:
@@
Figure 9: Dependence of the coincidences-to-accidentals ratio (CAR) on average power of the pump pulses for varying amounts of SpRS.
@@
Figure 10: (a) Schematic of the experimental setup showing the integration of classical and quantum traffic over deployed fiber infrastructure with multiplexing and filtering components. (b) Two-photon interference fringes showing maintained entanglement quality with visibilities of 59–68% in the presence of coexisting classical traffic.
@@
7.2 Polarization Stabilization Using Machine Learning
@@
We are currently studying the use of machine learning to find correlations between quantum information signals and classical control signals separated in wavelength. In the past, limitations caused by polarization mode dispersion (PMD) and the corresponding differential group delay, $\tau$), have required that control signals be placed within $\approx 2$ nm of their quantum counterparts [95]. We used an attention-based machine learning model to look for higher-order correlations between distantly separated signals. We generate training data for the model with a polarization scrambler, consisting of a motorized polarization controller to vary the SOP of the two signals followed by a 25 km spool of SMF-28 fiber to consider the effects of PMD (see Fig. 11(a)). After demultiplexing the two signals using a wavelength division multiplexer (WDM), we send each signal to a Stokes polarimeter, which reports each SOP as three Stokes parameters, which can be considered Cartesian coordinates on the Poincaré sphere. Our model directly predicts the SOP of the data signal based on the SOP of the control signal. Our most recent results [96] were able to predict the SOP of a signal separated from the control by 5 nm with an average root mean square error of 0.072 in individual Stokes parameters, as demonstrated in Fig. 11(b).
Figure 11: (a) Experimental setup to collect SOP polarization data for machine learning. The polarization controller acts as a scrambler, and the fiber spool is used to add the effect of polarization mode dispersion. (b) Main: Measured S1 parameter and our model predictions. The model faithfully tracks the state of polarization with a root mean square error of 0.072. Inset: Measured SOP states on the Poincaré sphere, demonstrating that we sample all possible SOPs.
@@
7.3 Quantum Network Orchestrator Prototype
To coordinate heterogeneous quantum devices and protocols on ARQNET, we developed a modular orchestrator that instantiates the three-plane abstraction introduced in Section 6. The orchestrator adopts a microservice architecture where each physical device such as entangled photon sources, polarization analyzers, or time taggers is encapsulated as a gRPC-based software agent. This design enables the orchestrator to issue high-level commands while device-specific control logic is handled locally, allowing protocols such as two-photon interference or quantum state tomography to be executed automatically across distributed nodes. Fig. 12 illustrates the orchestrator architecture, showing the gRPC-based interaction between device agents and the central control service.
@@ 7.3 Quantum Network Orchestrator Prototype
The orchestrator also incorporates polarization drift compensation as a gRPC service, enabling autonomous calibration and stabilization routines that replace manual realignment. To our knowledge, this is among the first orchestrators to implement a plane-based abstraction in a deployed quantum network testbed.
@@
7.4 Summary of Accomplishments and Open Challenges
Integration efforts on ARQNET have demonstrated feasibility across multiple layers. Coexistence studies showed that entanglement can be maintained alongside classical traffic, supported by an empirical model of Raman scattering and validated experimentally over deployed fiber. Polarization stabilization experiments using machine learning demonstrated that quantum channels can be predicted from classical control channels, with the potential to reduce the number of wavelengths needed for robust operation under dynamic conditions. A prototype quantum network orchestrator was also developed, implementing the three-plane abstraction with serial drivers for device control and gRPC-based coordination.
Challenges remain in extending coexistence studies to more traffic patterns and longer spans, demonstrating polarization stability over extended timescales, and scaling the orchestrator to manage heterogeneous entanglement across superconducting, erbium, and ytterbium platforms. These integration efforts complement the error management strategies of Section 5 and the architectural analysis of Section 6, advancing InterQnet toward its goals of heterogeneous demonstrations (Achieve) and guiding scalable design principles (Scale).
8 Conclusion
@@
Through this combination of experimental integration and systems-level study, InterQnet has advanced toward building a full-stack heterogeneous quantum network (InterQnet-Achieve) and guiding the design of scalable architectures (InterQnet-Scale). This work establishes a foundation for future demonstrations such as teleportation and Bell inequality violations on ARQNET, as well as scaling qLDPC-based purification to larger networks. Together, these results move the field closer to the realization of robust, heterogeneous quantum networking at national scale.

### v2

@@ VI-A1 Simulating Optical Quantum Communications
Kraus operators corresponding to the relevant noise models (e.g., amplitude damping, amplification, and depolarization) are applied to the matrix product states as they evolve through the system. These Kraus operators are applied probabilistically, with probabilities determined by the unnormalized weights of each trajectory [55]. Computing these probabilities for every iteration of a Monte Carlo experiment is computationally expensive and inefficient, thus we developed a novel trajectory caching scheme. Since most quantum channels in communication systems have certain trajectories that occur far more frequently than others, the required calculations for these common trajectories are performed only once, after which the cached results are reused for the majority of Monte Carlo samples.
VI-A2 Simulating Physical Qubits and Their Emission
@@
VI-A3 Simulating a Heterogeneous Quantum Network
To help understand engineering challenges in integrating heterogeneous nodes in a quantum network, we have developed in [82] faithful models of Yb-atom and superconducting nodes, as well as QFC and quantum transducers. In this preliminary study, we implemented entanglement generation and swapping protocols for time-bin encoded photons that account for disparate clock rates, quantum frequency conversion, and transducer losses/noise introduced by the heterogeneity. Using extensive simulations, we mapped the rate–fidelity trade space and identified the dominant bottlenecks unique to heterogeneous systems.
Figure 8 shows the simulation results for entanglement generation between two superconducting nodes connected via a Yb-atom repeater. We vary the transmon’s $T_{1}$ coherence time from its default of 0.5ms0.5~ms to a maximum of 10ms10~ms. For the optimal parameters (blue lines in the plots), all efficiency parameters are set to $100\%$, and noise rates are $0\%$. In contrast, the default device parameters assume near-term optimistic values, such as a 50% collection efficiency for Yb and a 60% transducer efficiency (see Table I in [82]). As expected, the rate with optimal parameters ($\sim 7~Hz$) is higher than with the default ($\sim 2~Hz$). We observe that the fidelity remains close to 0 for default parameters, while an increase in coherence time leads to tangible increases in fidelity (up to >0.5>0.5 at $T_{1}=10ms$).
Figure 8: Simulation results for entanglement generation between two superconducting nodes connected via an Yb-atom repeater, reproduced from [82]. We vary the coherence time of the superconducting node and compare the entanglement rate (a) and fidelity (b) for both currently achievable and optimal device parameters in SeQUeNCe.
VI-B Entanglement witnessing in the logical plane of quantum networks
@@
VI-C Architectural Approaches to the Control and Service Plane of Scalable Quantum Networks
@@
Here we study the trade-off of generating entanglement on demand vs. continuously generating link-level entanglement, with the goal of reducing the time to serve user requests. In prior work [61] we proposed a theoretical framework that enables the continuous generation of entangled pairs between neighboring nodes, similar to [118] and [46]. This approach can adapt to incoming client requests to reduce the time to serve future requests. In [61], however, the theoretical framework is evaluated only through small-scale, ad hoc, and naïve time-slotted simulations. The challenge is to implement the theoretical framework into a comprehensive quantum network protocol and simulate the protocol in a discrete-event quantum network simulator at a large scale, including realistic physical parameters.
@@
The challenges were addressed in [131], where we designed and implemented the Adaptive, Continuous entanglement generation Protocol (ACP). ACP can reduce the time to serve a client request by letting a node attempt to generate entangled pairs with neighboring nodes in the background continuously, regardless of incoming requests. The selection of neighbor nodes adapts to the request traffic by a feedback control algorithm. Additionally, entanglement distillation is used to enhance the fidelity of entangled pairs. We extended SeQUeNCe to support the simulation of ACP. In particular, we added a single-heralded entanglement generation protocol that accounts for pregenerated entangled pairs.
Figure 9: Simulation results for the 200-node autonomous system topology network, plots reproduced from [131]. (a) ACP outperforms on-demand-only (ODO) and uniform-continuous protocol (UCP) policies by having a 50+% lower time to serve than UCP. (b) ACP enhanced with entanglement distillation can increase the fidelity of the end-to-end entangled pairs by as much as 0.05 compared with UCP and 0.04 compared with ODO, both without entanglement distillation. Both plots show the adaptiveness as the traffic pattern changes in the middle of the simulation.
In [131], we showed promising results when we evaluated the ACP in the SeQUeNCe simulator under various scales (2-node, 20-node, and 200-node networks) and realistic physical parameters. Figure 9 shows selected results from our prior work. Although we assumed homogeneous quantum memories, ACP can reduce the time to serve by up to 94% and improve fidelity by 0.05 in large-scale simulations. We are now working on integrating the models described in Section VI-A3 and ACP to simulate large-scale heterogeneous quantum networks.
VI-D Summary of Accomplishments and Open Challenges
We have made significant progress in creating faithful and computationally efficient simulations of the infrastructure plane of quantum networks. We developed simulation tools based on tensor networks to simulate optical communications in multidimensional bosonic states efficiently. These tools will serve as the foundation to evaluate the qLDPC entanglement distillation protocols described in Section V-D. Our realistic implementation of an adaptive strategy to continuously generate link-level entanglement significantly outperforms the on-demand-only strategy by reducing the time to serve by up to 94% and by improving fidelity by 0.05 in large-scale simulations. We are now scaling and optimizing this for hybrid links, larger networks, and centralized vs. decentralized control paradigms. Collectively, these efforts build a scalable, realistic foundation for deploying and evaluating next-generation quantum networks.
Future work includes integrating devices and error correction models into SeQUeNCe and comparing centralized vs. distributed control schemes in the control and service plane of quantum networks. We will conduct studies on hybrid quantum error management strategies similar to [94]. While the authors focused on linear repeater chains of trapped ions with eight hops maximum, we will scale our simulations to complex topologies considering hundreds of heterogeneous nodes and the protocols developed in Section V. As described in [8], we are also interested in investigating under what conditions services level agreements between users and the network are required and what behaviors can be observed under different traffic patterns and network topologies.
VII Integration
@@
Integration efforts highlight the co-design aspects of the InterQnet project (see Fig. 1). Our current work focuses on the coexistence of quantum and classical signals, stabilization of quantum channels under realistic fiber conditions, and orchestration of heterogeneous devices. These activities are essential to InterQnet-Achieve, where a three-node heterogeneous network will demonstrate entanglement distribution across superconducting, erbium, and ytterbium platforms, to address challenges C1 and C2. Moreover, results of these integration activities inform InterQnet-Scale by identifying key constraints and performance trade-offs for realistic architectures (see challenge C3), improving the realism of our models and increasing confidence in our simulations.
Figure 10: Dark fiber optics network in the great Chicago area, including Argonne and partner institutions. The inset shows Argonne’s quantum network testbed (ARQNET), where InterQnet integration experiments are conducted.
VII-A Quantum-Classical Coexistence over Deployed Fiber
@@
We prototyped an on-demand entanglement distribution service over ARQNET (see inset in Fig. 10), addressing the challenge of spontaneous Raman scattering (SpRS) induced by coexisting classical signals [56]. We developed an empirical model that quantitatively describes the relationship between the coincidences-to-accidentals ratio (CAR) and pump power of entangled photon sources in the presence of classical traffic:
@@
Figure 11: Dependence of the coincidences-to-accidentals ratio (CAR) on average power of the pump pulses for varying amounts of SpRS. Plot reproduced from [56].
@@
Figure 12: Schematic of the experimental setup in [56], showing the integration of classical and quantum traffic over deployed fiber infrastructure with multiplexing and filtering components.
@@
Figure 13: Two-photon interference fringes showing maintained entanglement quality with visibilities of 59–68% in the presence of coexisting classical traffic. Reproduced from [56].
@@
VII-B Polarization Stabilization Using Machine Learning
@@
Figure 14: Experimental setup used in [39] to collect SOP polarization data for machine learning. The polarization controller acts as a scrambler, and the fiber spool is used to add the effect of polarization mode dispersion.
We are currently studying the use of machine learning to find correlations between quantum information signals and classical control signals separated in wavelength. In the past, limitations caused by polarization mode dispersion (PMD) and the corresponding differential group delay ($\tau$), have required that control signals be placed within $\approx 2$ nm of their quantum counterparts [128]. We used an attention-based machine learning model to look for higher-order correlations between distantly separated signals. We generate training data for the model with a polarization scrambler, consisting of a motorized polarization controller to vary the SOP of the two signals followed by a 25 km spool of SMF-28 fiber to consider the effects of PMD (see Fig. 14). After demultiplexing the two signals using a wavelength division multiplexer (WDM), we send each signal to a Stokes polarimeter, which reports each SOP as three Stokes parameters, which can be considered Cartesian coordinates on the Poincaré sphere. Our model directly predicts the SOP of the data signal based on the SOP of the control signal. Our most recent results [39] were able to predict the SOP of a signal separated from the control by 5 nm with an average root mean square error of 0.072 in individual Stokes parameters, as demonstrated in Fig. 15.
Figure 15: Main: Measured S1 parameter and our model predictions. The model faithfully tracks the state of polarization with a root mean square error of 0.072. Inset: Measured SOP states on the Poincaré sphere, demonstrating that we sample all possible SOPs. Reproduced from [39].
@@
VII-C Quantum Network Orchestrator Prototype
To coordinate heterogeneous quantum devices and protocols on ARQNET, we developed a modular orchestrator that instantiates the three-plane abstraction introduced in Section VI. The orchestrator adopts a microservice architecture where each physical device (such as entangled photon sources, polarization analyzers, or time taggers) is encapsulated as a gRPC-based software agent. This design enables the orchestrator to issue high-level commands while device-specific control logic is handled locally, allowing protocols such as two-photon interference or quantum state tomography to be executed automatically across distributed nodes. Fig. 16 illustrates the orchestrator architecture, showing the gRPC-based interaction between device agents and the central control service.
@@ VII-C Quantum Network Orchestrator Prototype
The orchestrator also incorporates polarization drift compensation as a gRPC service, enabling autonomous calibration and stabilization routines that replace manual realignment. Although our orchestrator follows the three-plane abstraction, the implementation is rooted in centralized control and SDN design principles. This approach aligns with current efforts from QUANT-NET [130] and the Quantum Internet Alliance [9].
VII-D Summary of Accomplishments and Open Challenges
Current integration efforts on ARQNET lay the foundation for the final InterQnet-Achieve demonstration. The ARQNET topology imposes a real constraint on our design because fiber links connecting the laboratories hosting heterogeneous devices and the integrated BSM and QFC have only two fiber strands; thus the coexistence of quantum and classical traffic will be unavoidable. Coexistence studies showed that entanglement can be maintained alongside classical traffic, supported by an empirical model of Raman scattering and validated experimentally over deployed fiber. Polarization stabilization experiments using machine learning demonstrated that quantum channels can be predicted from classical control channels, with the potential to reduce the number of wavelengths needed for robust operation under dynamic conditions. While the final heterogeneous devices in our testbed will rely on time-bin encoded photons, having polarization stabilization can improve entanglement rates. A prototype quantum network orchestrator was also developed, implementing centralized control with serial drivers for device control and gRPC-based coordination.
Challenges remain in extending coexistence studies to more traffic patterns and longer spans and demonstrating polarization stability over extended timescales. Data collected in ARQNET experiments can be used to validate SeQUeNCe models, as we have already shown in [112]. Scaling the orchestrator to manage heterogeneous entanglement across superconducting, erbium, and ytterbium platforms will require continuous development and integration of the software framework. Our current prototype provided valuable information on the engineering challenges of deploying a centralized controller and agents distributed across an institutional LAN composed of multiple security domains (i.e., firewalls). More work is required on improving our abstractions toward integration of the orchestrator with the local control systems of our heterogeneous qubit platforms.
[…]
@@ VII-D Summary of Accomplishments and Open Challenges
VIII Conclusion
@@
Through this combination of experimental integration and systems-level study, InterQnet has advanced toward building a full-stack heterogeneous quantum network (InterQnet-Achieve) and guiding the design of scalable architectures (InterQnet-Scale). This work establishes a foundation for future demonstrations such as teleportation and Bell inequality violations on ARQNET, as well as scaling qLDPC-based purification to larger networks. Together, these results move the field closer to the realization of robust, heterogeneous quantum networking at a national scale.